OpenStudy (bakablinkie):

@TheSmartOne -2.352 as a fraction

OpenStudy (evoker):

To turn a non repeating decimal into a fraction, take the number without the decimal as the numerator, and the place of the last digit 10th, Hundreth and so on for the denominator.

TheSmartOne (thesmartone):

Is 352 repeating? If it is, follow this example Let's say x = 0.111 (repeating) Multiply both sides by 10 10x = 1.11 Now subtract x from 10x 10x - x = 1.11 - 0.11 9x = 1 x = 1/9 so 0.111 repeating = 1/9

OpenStudy (unklerhaukus):

-2.352 = -2.352/1 = -2352/1000 = -1176/500 = -588/250 = -294/125 = -294/125 = -1 169/125 = -2 44/125